The vapor pressure of water is 23.76 mm Hg at 25 °C. A nonvolatile, nonelectrolyte that dissolves in water is urea.

Calculate the vapor pressure of the solution at 25 °C when 6.846 grams of urea, CH4N2O (60.10 g/mol), are dissolved in 207.3 grams of water.

water = H2O = 18.02 g/mol.

VP(solution) = mm Hg
Answer #1

Using, (P°-P)/P° = x

Where P° = vapour pressure of water = 23.76 mmHg

P = vapour pressure of solution

Moles of urea = mass/molar mass

= 6.846/60.10 = 0.1139 moles

Moles of water = 207.3/18 = 11.517 moles

Mole fraction of urea = moles of urea/total moles

= 0.1139/(0.1139+11.517)

= 0.1139/11.63 = 0.00979

Substituting values in equation,

(23.76-P)/23.76 = 0.00979

23.76 - P = 23.76×0.00979

23.76-P = 0.2327

P = 23.76-0.2327

= 23.53 mmHg

Please give feedback!